# Dynamic Prototype Rehearsal for Continual Learning in ECG Arrhythmia Detection

### v1

@@ Abstract
Index Terms—  ECG, heartbeat classification, Continual learning
1 Introduction
@@ 2 Related Work
Fig. 1: The overall framework of our proposed method.
3 Proposed Method
@@
Table 1: The performance of our method across three scenarios of time, class and lead incremental compared to prior works.
@@
Mean ()($\uparrow$)( ↑ ) AvgAUC ()($\uparrow$)( ↑ ) BWT ()($\uparrow$)( ↑ ) Mean ()($\uparrow$)( ↑ ) AvgAUC ()($\uparrow$)( ↑ ) BWT ()($\uparrow$)( ↑ ) Mean ()($\uparrow$)( ↑ ) AvgAUC ()($\uparrow$)( ↑ ) BWT ()($\uparrow$)( ↑ )
@@
Fig. 2: Performance during training on different sessions.
Fig. 3: Impact of no. clusters.
4 Experiments
@@
Table 2: Ablation study.
@@
Table 3: The effect of rehearsal memory size.
@@
Results. In Table 1 we present the performance of our model in comparison to state-of-the-art methods on three CL setups, namely time-incremental, class-incremental, and lead-incremental learning. We utilize two different rehearsal memory sizes for our method, small ($r_{\mathcal{M}}=0.25$) and large ($r_{\mathcal{M}}=0.75$), where $r_{\mathcal{M}}=|\mathcal{M}|/|\mathcal{D}$—. We observe that our model outperforms prior works, setting a new state-of-the-art. Among the competing methods, CLOPS demonstrates the overall second-best performance. Next, for a more detailed view of the performance of DREAM-CL, we report the AvgAUC across each session in Fig. 3. Here, we observe that our method consistently exhibits strong performances throughout the incremental sessions. By studying the type of incremental sessions used to evaluate our method, we observe that lead-incremental seems more challenging for CL as evidenced by the fluctuations, which could be due to the challenging nature of adapting to different random sets of ECG leads.
In Table 3, we explore the impact of the size of the rehearsal memory buffer on our method, where we observe that the memory size does not have a strong impact on the performance of our method, indicating the effectiveness of our sample selection strategy. Finally, we perform a sensitivity study on the number of clusters used to identify the key samples for the rehearsal memory and present the results (class-incremental scenario) in Fig. 3. In this figure we do not observe a dominant trend beyond a the first few clusters. In our experiments, we set $k=5$ for all setups.
5 Conclusion
@@
Acknowledgment. We thank NSERC, Mitacs, and Ingenuity Labs Research Institute for partially funding this research.

### v2

# Dynamic Prototype Rehearsal for Continual ECG Arrhythmia Detection

@@ Abstract
I Introduction
@@ II Related Work
Figure 1: The overall framework of our proposed method.
III Proposed Method
@@
TABLE I: The performance of our method across three scenarios of time, class and lead incremental compared to prior works.
@@
Mean $(\uparrow)$ AvgAUC $(\uparrow)$ BWT $(\uparrow)$ Mean $(\uparrow)$ AvgAUC $(\uparrow)$ BWT $(\uparrow)$ Mean $(\uparrow)$ AvgAUC $(\uparrow)$ BWT $(\uparrow)$
@@
Figure 2: Performance during training on different sessions.
IV Experiments
@@
TABLE II: Ablation study.
@@
Results. In Table I we present the performance of our model in comparison to state-of-the-art methods on three CL setups, namely time-incremental, class-incremental, and lead-incremental learning. We utilize two different rehearsal memory sizes for our method, small ($r_{\mathcal{M}}=0.25$) and large ($r_{\mathcal{M}}=0.75$), where $r_{\mathcal{M}}=|\mathcal{M}|/|\mathcal{D}$—. We observe that our model outperforms prior works, setting a new state-of-the-art. Among the competing methods, CLOPS demonstrates the overall second-best performance. Next, for a more detailed view of the performance of DREAM-CL, we report the AvgAUC across each session in Fig. 2. Here, we observe that our method consistently exhibits strong performances throughout the incremental sessions. By studying the type of incremental sessions used to evaluate our method, we observe that lead-incremental seems more challenging for CL as evidenced by the fluctuations, which could be due to the challenging nature of adapting to different random sets of ECG leads.
TABLE III: The effect of rehearsal memory size.
@@
In Table III, we explore the impact of the size of the rehearsal memory buffer on our method, where we observe that the memory size does not have a strong impact on the performance of our method, indicating the effectiveness of our sample selection strategy. Finally, we perform a sensitivity study on the number of clusters used to identify the key samples for the rehearsal memory and present the results (class-incremental scenario) in Fig. 3. In this figure we do not observe a dominant trend beyond a the first few clusters. In our experiments, we set $k=5$ for all setups.
Figure 3: Impact of no. clusters.
V Conclusion
@@
Acknowledgment
We thank NSERC, Mitacs, and Ingenuity Labs Research Institute for partially funding this research.